# Flexible Tails for Normalizing Flows

### v1

@@ 5 Conclusion
Most current methods for modelling extremes with normalizing flows use heavy tailed base distributions. We demonstrate this can perform poorly, due to extreme inputs to neural networks causing slow convergence. We present an alternative: use a Gaussian base distribution and a final transformation which can induce heavy tails. We propose the TTF transformation, equation (3), and prove it can indeed produce heavy tails. In our examples, a version of our method, TTF fix, has the best performance overall. For examples with heavy tails it is usually the best performer, or if not it is within a standard error of it. It does not damage performance for examples without heavy tails. Here all types of NF we considered perform well.
More generally, the experimental results show that final transformation methods generally outperform heavy base methods. We also compare two stage and one stage methods. The former begin by estimating the tail parameters and then train the rest of the flow. TTF fix is an example of this. The latter train all flow parameters jointly. We find that two stage methods often outperform one stage methods, although in other examples there is no difference. Interestingly, another two stage final transformation method – COMET flows (McDonald et al., 2022)—often performs competitively, which we discuss further in Appendix F.
@@ 5.1 Limitations and Future Work
Another potential limitation is that the TTF transformation affects both the body and tail of the output distribution. The preceding layers of the flow must learn to model the body of the distribution, and also adapt to the final layer. It would be appealing to decouple the transformations, so that they can concentrate separately on the body and tail, which would hopefully make optimisation easier. This could be achieved by somehow ensuring the final transformation is approximately the identity in the body region.
Our final transformation converts each margin to a heavy tailed distribution. This is limiting, as the choice of directions to transform depends on an arbitrary choice of parameterisation. Future work could try to address this by using an autoregressive structure (Papamakarios et al., 2021) for the tail parameters for each marginal, or by adding a linear layer (e.g. that of Oliva et al., 2018) after our transformation.
Possible future work is to design methods which incorporate more extreme value theory properties. For instance, our method doesn’t explicitly allow for tail dependence (Coles et al., 1999). This would be an interesting direction for future work, for instance by adapting the manifold copula method proposed by McDonald et al. (2022). Also, it would be interesting to design multivariate transformations with the max-stable property (Coles, 2001).
As mentioned above, our best experimental results are for a two stage method. A corresponding method for VI would be useful, such as using the static analysis method of Liang et al. (2024) to generate initial tail weight estimates. Also, it would be desirable to understand why a one stage method has worse performance, and try to improve this.
In this paper we’ve only included a proof-of-concept VI example on an artificial target. In exploratory work, we found it difficult to improve VI results on real examples using our method. One reason is that our experiments only show major improvements for sufficiently heavy tails (e.g. $\nu<2$ in Tables 2, 3), which we found to be rare for real posteriors. The limitations of our transformation outlined above may also contribute.
Finally, a potential future application is to simulation based inference. Here neural likelihood estimation methods (Papamakarios et al., 2019) use density estimation with NFs to estimate a likelihood function from simulated data. Our approach could improve results when the simulated data is heavy tailed. More generally, our transformation may be useful to produce heavy tails with other generative models.
@@ Acknowledgements
Thanks to Miguel de Carvalho, Seth Flaxman, Iain Murray, Ioannis Papastathopoulos, Scott Sisson, Jenny Wadsworth and Peng Zhong for helpful discussions. Tennessee Hickling is supported by a PhD studentship from the EPSRC Centre for Doctoral Training in Computational Statistics and Data Science (COMPASS). A preliminary version of this work (Hickling and Prangle, 2023), using a similar transformation was presented to the MIDAS workshop at ECML-PKDD 2023.
Appendix A TTF details
@@
$\lambda_{s}=\begin{cases}\lambda_{+}&\text{for }s=1,\\ \lambda_{-}&\text{for }s=-1.\end{cases}$
First, Section A.1 reviews some background on the complementary error function. Then Section A.2 presents a motivation for the transformation, and Section A.3 discusses some of its properties.
A.1 Complementary error function
@@
$F_{N}(z)=\frac{1}{2}(1+\operatorname{erf}[z/\sqrt{2}]).$ (5)
@@
Efficient numerical evaluation of $\operatorname{erfc}$ and its gradient is possible, as it is a standard special function (Temme, 2010), implemented in many computer packages. For instance PyTorch provides the torch.special.erfc function. Later we also require $\operatorname{erfc}^{-1}$, which is less commonly implemented directly. However using (6) we can compute $\operatorname{erfc}^{-1}$ using the more common standard normal quantile function $F_{N}^{-1}$. For instance PyTorch implements this as torch.special.ndtri (although this can have problems for small inputs—see Appendix D.3).
A.2 Motivation
@@
Here $\lambda_{+}>0,\lambda_{-}>0$ are shape parameters for the positive and negative tails. We use the notation $s=\operatorname{sign}(u)$, with $\lambda_{s}$ defined as earlier.
@@
so there is a standard special function to compute $1-|u|$ directly. Large $|z|$ gives $1-|u|\approx 0$, which can be represented to high accuracy in floating point arithmetic, avoiding the catastrophic rounding which could result from working directly with $u$.
@@
A.3 Properties
Here we derive several properties of the TTF transformation, $R$. These include expressions which are useful in an automatic differentiation implementation of $R$ for normalizing flows: $R^{-1}$ and derivatives of $R$ and $R^{-1}$.
@@ A.3 Properties
Appendix B Asymptotic results
This section proves asymptotic results on the tails produced by COMET flows and our TTF transformation. Our main aim to to prove whether or not the distributions are in the Fréchet domain of attraction $\Theta_{\lambda}$ (see Definition 3). We will also find the tail behaviour in detail for some special cases.
Throughout this section we use the following asymptotic notation. In particular, $\sim$ is not used to describe probability distributions, as it is elsewhere in the paper.
@@ Appendix B Asymptotic results
B.1 Background
The material in this section will be used in the proofs later in this appendix. The first results are on regularly varying functions, and based on the presentation in Nair et al. (2022).
@@ B.1 Background
The following result is Lemma 2.7 of Nair et al.
Theorem 6
@@
The following result is Theorem 2.8 of Nair et al.
Theorem 7
@@
Of interest is a Gaussian base distribution transformed by a Lipschitz transformation, such as most standard normalizing flows. The result shows that the output is in $\overline{\mathcal{E}^{2}}$, and not the Fréchet domain of attraction.
B.2 COMET flow tails
@@ B.3 TTF tails
    For $\mu=0,\sigma=1$, $X$ has density $q(x)\sim q_{GPD}(x;\lambda,2^{1/(2+1/\lambda)})$, a GPD density with shape parameter $\lambda$ and scale $2^{1/(2+1/\lambda)}$ (see Definition 13 below.)
@@ A GPD with shape $\lambda>0$, location 0, and scale $\sigma>0$ has density
Proof  Let $F_{N}(\cdot)$ be the $\mathcal{N}(0,1)$ cdf. The result is a corollary of the following asymptotic expansion from Fung and Seneta (2018) (for $p\to 0$)
@@ B.3.2 Proof of Theorem 12
Appendix C TTF Transform: Variations and Related Work
@@
An intuitively appealing alternative transformation which converts Gaussian tails to heavy tails is
@@
where $F_{T}$ and $F_{N}$ are cdfs of Student’s $T$ and Gaussian distributions (both with location zero and scale one). Rather than use the Student’s $T$ distribution for the base distribution, as in the prior work reviewed in Section 1.1.2, this moves the use of the $T$ distribution to the final transformation. Unfortunately, $R_{\text{cdf}}$ is not practical as $F^{-1}_{T}$ does not have a closed form allowing automatic differentiation.
@@
Another related result is from Troshin (2022). This paper investigates necessary and sufficient conditions for transformations to map $\mathcal{N}(0,1)$ inputs to the Fréchet domain of attraction. This produces a family of functions which includes $R$ and $R_{\text{alt}}$—see Troshin’s Remark 2. However, unlike our Appendix B, this work does not provide results for $\mathcal{N}(\mu,\sigma^{2})$ inputs.
Appendix D Implementation Details
@@ D.1 Normalizing Flows
Our NF layers (autoregressive rational quadratic spline layer, and autoregressive affine layers) output the required transformation parameters from masked neural networks. These use 2 hidden layers, each with a width equal to the input dimension plus 10.
D.2 Sampling Student’s T
Let $t_{\nu}$ represent a Student’s $T$ distribution with $\nu$ degrees of freedom. Algorithm 1 draws samples, while allowing the reparameterization trick i.e. differentiation with respect to $\nu$.

### v2

@@ 5 Conclusion
Most current methods for modelling extremes with NFs use heavy tailed base distributions. We demonstrate this can perform poorly, due to extreme inputs to neural networks causing slow convergence. We present an alternative: use a Gaussian base distribution and a final transformation which can induce heavy tails. We propose the TTF transformation, equation (3), and prove it can indeed produce heavy tails. For density estimation, our TTF methods outperform existing methods on real and synthetic data. Jointly learning the tail parameters and the rest of the flow is usually the best approach. However for variational inference, fixing the tail parameters is beneficial, perhaps by improving optimisation stability. Our approaches do not damage performance in examples without heavy tails – here all types of NF we considered perform well.
@@ 5.1 Limitations and Future Work
Another potential limitation is that the TTF transformation affects both the body and tail of the output distribution. The preceding layers of the flow must learn to model the body of the distribution, and also adapt to the final layer. It would be appealing to decouple the transformations, so that they can concentrate separately on the body and tail, which could make optimisation easier. This could be achieved by somehow ensuring the final transformation is approximately the identity in the body region.
Possible future work is to design methods which incorporate more extreme value theory properties. For instance, our method doesn’t explicitly allow tail dependence (Coles et al., 1999). This would be an interesting future direction e.g. by adapting the manifold copula method of McDonald et al. (2022). Also, it would be interesting to design multivariate transformations with the max-stable property (Coles, 2001).
We found that fixing the tail parameters to known values performed well in our VI example on an artificial target. However in real VI applications, the tail parameters aren’t known, motivating methods to estimate them e.g. using static analysis of probabilistic programs (Liang et al., 2024). Also, in exploratory work on real VI applications, we found it difficult to improve VI results using our method. A possible reason is that our experiments only show major improvements for especially heavy tails ($\nu\leq 1$ in Table 3), which we found to be uncommon for real posteriors.
A potential future application is simulation based inference. Here neural likelihood estimation methods (Papamakarios et al., 2019) use NF density estimation to estimate a likelihood function from simulated data. Our approach could improve results when the simulated data is heavy tailed.
Finally, it would be interesting to use our tail transformation with other generative models. Here we briefly discuss differential equation based methods e.g. continuous normalizing flows, diffusion models, flow matching (Lipman et al., 2023). These sample $x(0)$ from a base distribution such as $\mathcal{N}(0,I)$, apply a differential equation $\frac{dx}{dt}=u(x,t)$ (or an SDE also including a diffusion term), and output $x(1)$. The vector field $u$ is a neural network, trained to produce a desired output distribution. Our work suggests it would be difficult to train a vector field for heavy tailed targets. To see this, suppose we desire output $x(1)=x^{*}$. By continuity, we need $x(1-\delta)\approx x^{*}$ for small $\delta$. So we must evaluate $u$ for input similar to $x^{*}$, which can be extreme for a heavy tailed target. We’ve argued neural networks are hard to train with extreme inputs. This motivates using our tail transform as a final transformation to $x(1)$, effectively lightening the tails of the target, avoiding extreme $x^{*}$ values. Either joint training or a two-stage approach could be tried, similar to TTF and TTFfix.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
Thanks to Miguel de Carvalho, Seth Flaxman, Iain Murray, Ioannis Papastathopoulos, Scott Sisson, Jenny Wadsworth, Peng Zhong and anonymous reviewers for helpful discussions. Tennessee Hickling is supported by a PhD studentship from the EPSRC Centre for Doctoral Training in Computational Statistics and Data Science (COMPASS). A preliminary version of this work appeared as Hickling and Prangle (2023).
Appendix A Gaussian mixtures are light tailed
As suggested by a reviewer, here we prove that Gaussian mixtures are light tailed. Then, by Theorem 2, using a Gaussian mixture base distribution with a Lipschitz NF produces light tailed output. Amiri et al. (2022) suggest this combination (reviewed in Section 1.1.3), and we consider it in our experiments under the name m_normal.
Suppose the real valued random variable $Z$ has density
$q(z)=\sum_{i=1}^{m}w_{i}q_{i}(z),$
where $\sum_{i=1}^{m}w_{i}=1$, $w_{i}\in(0,1)$ and $q_{i}(z)$ is a $Z_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$ density. The moment generating function of $Z_{i}$ is
$M_{Z_{i}}(\lambda)=\operatorname{\mathbb{E}}[\exp(\lambda Z_{i})]=\exp[\lambda\mu_{i}+\lambda^{2}\sigma_{i}^{2}/2].$
Hence the moment generating function of $Z$ is
$M_{Z}(\lambda)=\operatorname{\mathbb{E}}[\exp(\lambda Z)]=\sum_{i=1}^{m}w_{i}M_{Z_{i}}(\lambda)=\sum_{i=1}^{m}w_{i}\exp[\lambda\mu_{i}+\lambda^{2}\sigma_{i}^{2}/2].$
So $M_{Z}(\lambda)<\infty$ for all $\lambda>0$, and $Z$ is light tailed under Definition 1.
In this paper we concentrate on univariate properties. However, note that for $Z$ which is a mixture of multivariate Gaussians, we can consider any univariate random variable of the form $Z^{\prime}=a^{T}Z$ (with $a^{T}a\neq 0$). Since $Z^{\prime}$ is clearly a mixture of univariate Gaussians, it is light tailed by the argument above.
Appendix B TTF details
@@
$\lambda_{s}=\begin{cases}\lambda_{+}&\text{for }s=1,\\ \lambda_{-}&\text{for }s=-1.\end{cases}$ (5)
First, Section B.1 reviews some background on the complementary error function. Then Section B.2 presents a motivation for the transformation, and Section B.3 discusses some of its properties.
B.1 Complementary error function
@@
$F_{N}(z)=\frac{1}{2}(1+\operatorname{erf}[z/\sqrt{2}]),$ (6)
@@
Efficient numerical evaluation of $\operatorname{erfc}$ and its gradient is possible, as it is a standard special function (Temme, 2010), implemented in many computer packages. For instance PyTorch provides the torch.special.erfc function. Later we also require $\operatorname{erfc}^{-1}$, which is less commonly implemented directly. However using (7) we can compute $\operatorname{erfc}^{-1}$ using the more common standard normal quantile function $F_{N}^{-1}$. For instance PyTorch implements this as torch.special.ndtri (although this can have problems for small inputs—see Appendix G.5).
B.2 Motivation
@@
Here $\lambda_{+}>0,\lambda_{-}>0$ are shape parameters for the positive and negative tails. We use the notation $s=\operatorname{sign}(u)$, with $\lambda_{s}$ defined as in (5).
@@
so there is a standard special function to compute $1-|u|$ directly. Large $|z|$ gives $1-|u|\approx 0$, which can be represented to high accuracy in floating point arithmetic, avoiding catastrophic rounding which could result from working directly with $u$.
@@
B.3 Properties
Here we derive several properties of the TTF transformation, $R$. These include expressions which are useful in an automatic differentiation implementation of $R$ for NFs: $R^{-1}$ and derivatives of $R$ and $R^{-1}$.
@@ B.3 Properties
Appendix C Asymptotic results
This appendix proves asymptotic results on the tails produced by COMET flows and our TTF transformation. Our main aim to to prove whether or not the distributions are in the Fréchet domain of attraction $\Theta_{\lambda}$ (see Definition 3). We will also find the tail behaviour in detail for some special cases.
Throughout this appendix we use the following asymptotic notation. In particular, $\sim$ is not used to describe probability distributions, as it is elsewhere in the paper.
@@ Appendix C Asymptotic results
C.1 Background
The material in this section will be used in the proofs later in this appendix. The first results are on regularly varying functions, based on the presentation in Nair et al. (2022).
@@ C.1 Background
Theorem 6 (Nair et al., Lemma 2.7)
@@
Theorem 7 (Nair et al., Theorem 2.8)
@@
Of interest is a Gaussian base distribution transformed by a Lipschitz transformation, such as most standard NFs. The result shows that the output is in $\overline{\mathcal{E}^{2}}$, and not the Fréchet domain of attraction.
C.2 COMET flow tails
@@ C.3 TTF tails
    For $\mu=0,\sigma=1$, $X$ has density $q(x)\sim q_{GPD}(x;\lambda,2^{1/(2+1/\lambda)})$, a GPD density with shape parameter $\lambda$ and scale $2^{1/(2+1/\lambda)}$ (see Definition 13 below).
@@ A GPD with shape $\lambda>0$, location 0, and scale $\sigma>0$ has density
Proof  Let $F_{N}(\cdot)$ be the $\mathcal{N}(0,1)$ cdf. The result is a corollary of the following asymptotic expansion from Fung and Seneta (2018), for $p\to 0$
@@ C.3.2 Proof of Theorem 12
Appendix D TTF Transform: Variations and Related Work
@@
As well as those discussed below, we expect many other variations are possible with equivalent asymptotic properties, providing much scope for potential future study.
Student’s $T$ CDF Transform
A straightforward alternative transformation which converts Gaussian tails to heavy tails is
@@
where $F_{T}$ and $F_{N}$ are cdfs of Student’s $T$ and Gaussian distributions (both with location zero and scale one). Rather than use the Student’s $T$ distribution for the base distribution, as in the prior work reviewed in Section 1.1.2, this moves the use of the $T$ distribution to the final transformation. Unfortunately, $F^{-1}_{T}$ does not have a closed form allowing automatic differentiation, so it lacks the flexibilty of TTF. In particular we cannot learn the tail weight parameters. However, $R_{\text{cdf}}$ could be applied in a two step procedure for density estimation similar to TTFfix (see Section 3.4). Exploratory work suggests this has comparable performance to TTFfix.
Unit Scale Transformation
@@
Another related result is from Troshin (2022). This paper investigates necessary and sufficient conditions for transformations to map $\mathcal{N}(0,1)$ inputs to the Fréchet domain of attraction. This produces a family of functions which includes $R$ and $R_{\text{alt}}$—see Troshin’s Remark 2. However, unlike our Appendix C, this work does not provide results for $\mathcal{N}(\mu,\sigma^{2})$ inputs.
Appendix E Universality
Here we prove some universality properties relating to the TTF transformation, as referred to in Section 3.5, which also includes a broader discussion on universality.
E.1 Background
We define universality for our purposes in Sections E.1E.2 as follows.
Definition 16
Let $\mathcal{S}$ be a set of transformations $s:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Let $Z\sim\mathcal{N}(0,I)$ be the usual base distribution. We call $\mathcal{S}$ universal if for every random variable $X$ on $\mathbb{R}^{d}$ there is a sequence $s_{n}\in\mathcal{S}$ such that $s_{n}(Z)\to X$ in the sense of weak convergence as $n\to\infty$.
Universality has been proved for some NFs including SoS polynomial flows and Neural Autoregressive Flows (Huang et al., 2018; Jaini et al., 2019; Kobyzev et al., 2020).
For all these results the set of transformations $\mathcal{S}$ comprises NFs of a particular type with unbounded capacity. The sequence $s_{n}$ typically requires increasing capacity as $n\to\infty$. So Definition 16 means that an idealised sequence of NFs can approximate any target distribution. In practice, capacity is limited and only an imperfect approximation is usually achievable.
For many other common NFs it is unknown whether universality holds: see Jaini et al., 2019 Table 1, but note progress has been made on coupling flows, discussed in Section E.3.
E.2 TTF preserves universality
Let $\mathcal{S}$ be the set of transformations corresponding to a universal NF. Let $\mathcal{R}$ be the set of TTF transformations under all valid parameter choices. Let $\tilde{\mathcal{S}}=\{r\circ s:r\in\mathcal{R},s\in\mathcal{S}\}$. We now prove that universality of $\mathcal{S}$ implies universality of $\tilde{\mathcal{S}}$. Note that $\mathcal{R}$ does not contain the identity, so the result is not immediate.
Fix any $\tilde{r}\in\mathcal{R}$ and some target $X$. Since $\mathcal{S}$ is universal, it contains a sequence $\{s_{n}\}$ such that $s_{n}(Z)\to\tilde{r}^{-1}(X)$. By the continuous mapping theorem (see e.g. Van der Vaart, 2000, Theorem 2.3), the continuity of $\tilde{r}$ implies $\tilde{r}\circ s_{n}(Z)\to X$, as required.
E.3 Coupling flows
Draxler et al. (2024) investigate universality of coupling flows. They replace Definition 16 by an alternative convergence metric specialised to the problem. Given a target random variable $X$ and a normalising flow $s$ acting on base random variable $Z\sim\mathcal{N}(0,I)$, it considers the closeness of $s^{-1}(X)$ to $\mathcal{N}(0,I)$ (in terms of the improvement possible by using $s\circ t$ where $t$ is a single coupling flow layer). See Definition 5.1 of Draxler et al. (2024) for full details.
Their main result – Theorem 5.4 of Draxler et al. 2024 – applies to any $X$ with full support, a continuous density and finite first and second moments. They show there exists a sequence of coupling flows $\{s_{n}\}$ achieving a convergence metric tending to zero.
The assumption that $X$ has finite first and second moments excludes sufficiently heavy tailed distributions from this result. This improves on earlier work such as Teshima et al. (2020) which looks at weak convergence in a bounded subset of $\mathbb{R}^{d}$, and so does not consider tails at all.
It is plausible that our TTF transformation can preserve the result of Draxler et al. (2024) and extend it to heavier tailed targets. Given some target $X$, the idea is to select $\tilde{r}\in\mathcal{R}$ such that $X^{\prime}=\tilde{r}^{-1}(X)$ has finite first and second moments. That is, $\tilde{r}^{-1}$ acts to lighten tails of $X$ which are too heavy. Take the sequence $\{s_{n}\}$ from applying the result of Draxler et al. (2024) to $X^{\prime}$. Then for target $X$ the sequence $\{\tilde{r}\circ s_{n}\}$ produces a convergence metric tending to zero.
However further results are needed to prove when a suitable $\tilde{r}$ exists. We leave this for future work.
Appendix F Neural Network Regression with Extreme Inputs
In Section 1, we claimed that neural network optimisation can perform poorly under heavy tailed input. This forms part of the motivation for proposing our methods. In Section 4.1 we illustrated this claim empirically for an example involving NFs. Here we provide further support in a simpler supervised learning example.
Our experiment considers the following simple regression problem
$\{X_{i}\}_{i=1}^{d}\sim t_{\nu},\quad Y\sim\mathcal{N}(X_{d},1).$
The $d$-dimensional input is heavy tailed. The output equals one of the inputs plus Gaussian noise. The other inputs act as nuisance variables.
Experimental Details
We consider a number of tail weight $(\nu$) and dimension ($d$) combinations and generate $5000$ train, validation and test samples for each. Our models are simple 2 layer multi-layer perceptrons with $50$ nodes in each hidden layer. The models are optimised with Adam to minimise mean square error, selecting the model with smallest validation loss found during training. We consider both sigmoid and ReLU activations. This is to investigate whether the main issue is saturation of sigmoid activation functions. The experiment was repeated 5 times, each trial sampling a new set of data.
Results
Table 4: Neural network regression example results. Values are the median test MSE over 5 trials, with the maximum provided in brackets. (We use median as it’s more robust than mean to extreme outliers. Maximum illustrates the presence and magnitude of such outliers.)
$d$ $\nu$ Sigmoid activation ReLU activation
30.0 1.01 (1.02) 1.01 (1.04)
5 2.0 1.74 (6.19) 1.06 (1.08)
1.0 1.57e+03 (8.57e+06) 3.45 (1.93e+03)
30.0 1.02 (1.02) 1.02 (1.05)
10 2.0 2.04 (7.28) 1.03 (1.05)
1.0 4.13e+03 (1.09e+06) 8.41 (49.6)
30.0 1.01 (1.04) 1.08 (1.09)
50 2.0 3.09 (5.26) 1.27 (12.3)
1.0 1.03e+04 (1.73e+06) 55.8 (6.59e+03)
30.0 1.02 (1.07) 1.15 (1.2)
100 2.0 3.41 (4.33) 1.60 (1.74)
1.0 1.02e+05 (1.94e+05) 336 (1.2e+03)
Appendix G Implementation Details
@@ G.1 Normalizing Flows
With the exception of the CLIMDEX example, our spline and autoregressive affine NF layers output the required transformation parameters from masked neural networks. These use 2 hidden layers, each with a width equal to the input dimension plus 10. Spline layers are configured to use a bounding box of $[-2.5,2.5]$ with 5 bins.
For the more complicated CLIMDEX dataset, which requires more capacity, we reuse the architecture and tuning choices used in Laszkiewicz et al. (2022), as discussed in Section 4.2.
G.2 Alternative Base Distributions
Standard NFs use a Gaussian base distribution, and a Student’s $T$ base distribution is also popular for heavy tailed targets. In addition to these, we also tried two alternative base distributions suggested by Amiri et al. (2022) which we describe here.
The first is a Gaussian mixture model base distribution (m_normal) with 5 components (preliminary work found 10 or 15 components had no significant difference in performance). For each component we optimise the $d$ dimensional mean and diagonal covariance terms along with other parameters.
The second alternative base distribution is $d$ independent generalised normal distributions (g_normal). The generalised normal family does not have Pareto tails, so we cannot fix them to have the true tail weight. Instead, we optimise the shape parameter of each marginal.
G.3 Tail Parameter Initialisation
For density estimation of heavy tailed data, the tail parameters must be initialised sufficiently high that large observations aren’t mapped to vanishingly low probability regions of the base distribution. If this does occur, it is possible to get numerical overflow during optimisation. For variational inference a related problem can occur: if some tail parameters are too high, then we can sample points of very low target density and get numerical overflow. In practice, we find that initialising $\lambda$ uniformly from $[0.05,1]$ provides sufficient stability.
For two-stage methods, Section 3.4 describes our initial stage of tail parameter estimation. Equivalent tail parameters are used for mTAF, TTFfix and COMET. In some cases, this estimation procedure finds that a marginal distribution is light tailed. In these cases, as mentioned in Section 3.3, we set the tail parameter to a small value, corresponding to a degree of freedom of $\nu=1000$.
G.4 Sampling Student’s T
Consider sampling from a Student’s $T$ distribution with $\nu$ degrees of freedom. Algorithm 1 implements this, while allowing the reparameterization trick i.e. differentiation with respect to $\nu$.